OpenStudy (anonymous):

How do you solve 49^x = sqrt7 ^(2x+6)?

OpenStudy (zzr0ck3r):

\[49^x = \sqrt{7^{2x+6}}\]?

OpenStudy (anonymous):

No, sorry, i meant 49^x = sqrt(7) ^ 2x+6

OpenStudy (anonymous):

\[\sqrt{7}^{2x+6}\] I guess

OpenStudy (anonymous):

49^x=7^x+3 ..by cancelling sqrt with power

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

zzr0ck3r how did you eliminate the square root on the 7 and put two there?

OpenStudy (zzr0ck3r):

\[\sqrt{7}=7^{\frac{1}{2}}\]

OpenStudy (anonymous):

aha

OpenStudy (zzr0ck3r):

\[49^x = \sqrt{7}^{2x+6}=(7^{\frac{1}{2}})^{2x+6}\\7^{2x}=7^{x+3}\\take \space \log_7\space of\space both\space sides\\2x=x+3\\x=3\]

OpenStudy (anonymous):

I understand it now... Thank you so much!!

OpenStudy (zzr0ck3r):

np